1/3+2/3 divided by 1/2x2/3
How would you solve this? Thank you!

Found 2 solutions by lynnlo, rfer:
Answer by lynnlo(4176)
Answer by rfer(16322)

(3/3)/(1/2)(2/3)
3/3*2/1*2/3=6/9=2/3